OpenStudy (babyslapmafro):

Please help me find the anti-derivative of 1/(4+x^2)

OpenStudy (babyslapmafro):

I'm trying to turn the 4 in the denominator into a 1 so that I may use arctan as the anti-derivative

OpenStudy (hba):

Integrate.

OpenStudy (babyslapmafro):

yes i am integrating

OpenStudy (babyslapmafro):

I'm actually finding the volume of the solid but I just need help finding the anti-derivative of this function

OpenStudy (hba):

Hmm.

OpenStudy (anonymous):

You can factorize out the 4 to get \[\frac{ 1 }{ 4 }\int\limits_{}^{}\frac{ 1 }{ \frac{ x^2 }{ 4 }+1 } \]

OpenStudy (anonymous):

Then use substituition by letting u=x/2